Intensity I(λ) = C − De^{−(λ−λ_{0})^{2}/2σ^{2}}

σ = \frac{FWHM}{2.35}

χ^{2} ≡ ∑_{i} \frac{(model_{i} − data_{i})^{2}}{(uncertainty_{i})^{2}}

dof ≡ (number of measurements ) − (number of parameters in fit)

reduced χ^{2} = \frac{χ^{2}}{dof}